Two rods of equal length (0.45 cm) and diameter (1.5 cm) are placed end to end. One rod is aluminum, the other is brass. If a compressive force of 8400 N is applied to the rods, how much does their combined length decrease?

1
Expert's answer
2022-04-11T12:08:09-0400

"\delta=\Sigma\frac{Fl_i}{E_iS}=\frac{4Fl}S(\frac 1{E_1}+\frac 1{E_2})=0.48~mm."